Shift Summary for Day Nov. 7, 1999

Hogan Nguyen (hogann@fnal.gov)
Sun, 7 Nov 1999 16:28:37 -0600 (CST)

Shift Summary for DAY shift, 11/07/99 (Sun)

On Shift: Rick Tesarek, Hogan

Hours of beam available at all intensities: 8h
Hours of beam available at nominal intensity: 8h
Hours of E799 data (special runs not included): 8h
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity: run stoppage

========================================================================
Runs ended this shift:

14919
14920

========================================================================
Accesses during shift:

None.

========================================================================
Overview of shift:

Beam available for the entire shift except for a few aborts.

Spill is smooth for the majority of the shift. We had occasional bouts of 2hz
spill structure. When this happened, we called MCR, who was able to fix
it.

Very few Ktevana errors.

========================================================================
Description of Shift:

Run Spill Time Note
--- ----- ---- ----
14919 07:00 Take over from owl shift. Good beam intensity
and structure. Spill timing is 40 seconds onspill,
45 seconds offspill.

14919 9:04 The current run is very long, making it hard to spot
new problems. Since the spill structure is bad, we
decide to start another run to get new histograms as
MCR is fixing our beam structure.

14920 9:10 New run started.

10:50 Mount split tape UPK014

Veto Fractions for RC6,7,10,SA's,HA are higher
than reference plots. Found by doing trg#veto.
SET notified. RJT suspects LRS429 DC offsets
drifted. Also, this might be due to pedestal
drifts.

RJT checks the effect of HA veto on trigger rates.
See logbook XXII page 182-3.

YTF errors. Could this be relaled to the
KQ readout problems.

14920 13:14 Run ends as the tape fills up.

Jason Hamm adds 1/2 nsec decay to RF going to Etot
to improve HCC threshold distribution

Update Fera Pedestal as per Valeri's request

14922 13:33 New BEAM_E799 run starts. Had to reload HCC for
the first few spills. Data good after spill 4.

Fera pedestal update helps the HA veto fractions
somewhat. Could still be a problem with LRS429 DC
offset.

========================================================================
Follow up Notes on Previous Shift

Still no B16 triggers.

========================================================================
Items to Follow up on Next Shifts

If at least 1hour down time, make access to check LRS429 offsets for the
HA and photon vetos.